Ayayai Magazine sold 11,280 annual subscriptions on August 1, 2020, for $20 each. Prepare Ayayai’s August 1, 2020, journal entry and the December 31, 2020, annual adjusting entry, assuming the magazines are published and delivered monthly. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Record journal entries in the order presented in the problem.)

Date Account Titles and Explanation Debit Credit

choose a transaction date

8/1/20 12/31/20 enter an account title enter a debit amount enter a credit amount

enter an account title enter a debit amount enter a credit amount

choose a transaction date

8/1/20 12/31/20 enter an account title enter a debit amount enter a credit amount

enter an account title enter a debit amount enter a credit amount
Answer #1

Answer:

Journal Entries:

Date Accounts title and Explanation Debit ($) Credit ($) Calculation
8/1/2020 Cash A/c Dr 225,600 = 11,280 * 20
To Unearned Subscription Revenue A/c 225,600
12/31/2020 Unearned Subscription Revenue A/c Dr 94,000 = 225,600 / 12 * 5 months
To Subscription Revenue A/c 94,000